OpenStudy (anonymous):

A string is 5m long under the tension of 2N. A transversal wave propagates in the string at speed of 40m/s. The mass of the string is? A. 6.25 gram B. 6.5 gram C. 6.75 gram D. 6.885 gram E. 6.90 gram

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

speed = square root (Tension / mass per unit length) \[s = \sqrt{\frac{ T }{ \mu }} \] \[\mu = T/s^2\] Mass of the string = mu * length of the string Does this help ?

OpenStudy (anonymous):

should just substitute the numbers into the equation @ShailKumar ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

mu means??

OpenStudy (anonymous):

\[\mu \] is mu i.e mass/length

OpenStudy (anonymous):

thank u i gt the answer A :) @ShailKumar

OpenStudy (anonymous):

Great! :)